OpenStudy (anonymous):

Graph the linear equation by solving y first -x+2y=4

OpenStudy (skullpatrol):

You are missing a very important word in the question ...

OpenStudy (anonymous):

solving for Y first

OpenStudy (skullpatrol):

Yes.

OpenStudy (skullpatrol):

Express y in terms of the other variable means write the equation in the form y = ...

OpenStudy (anonymous):

I have not done this stuff in over 12 year I do not understand how to start it off

OpenStudy (skullpatrol):

The plan is to use the idea of balance: whatever you do to one side you must do to the other to keep the balance.

OpenStudy (skullpatrol):

So, add x to both sides please.

OpenStudy (anonymous):

-x+2y=4x ?

OpenStudy (skullpatrol):

$$\huge -x+2y=4 $$ $$\huge -x+2y + x = 4 + x$$

OpenStudy (anonymous):

ok

OpenStudy (anonymous):